The sum of three numbers is 5. The first number minus the second plus the third is 1. The first minus the third is 3 more than the second. Find the numbers.
Answer by checkley77(12569)

x+y+z=5
x-y+z=1 multiply by -1 & add
-x+y-z=-1
x+y+z=5
------------
2y=4
y=4/2
y=2 ans.
x-z=y+3
x-y-z=3
x+y+z=5 add
----------
2x=8
x=8/2
x=4 ans.
4+2+z=5
6+z=5
z=5-6
z=-1 ans.